OpenStudy (anonymous):

Solve D = 1 over 4 (e – 3f) for e. You must show all work.

jhonyy9 (jhonyy9):

D=1/4(e-3f) multiply both sides with 4(e-3f) 4(e-3f)D=1(4(e-3f))/4(e-3f) after simplify you get 4(e-3f)D=1 4eD-12fD=1 4eD=1+12fD e=(1+12fD)/4D